# Predictable Compression Failures Why Language Models Actually Hallucinate

### v1

@@ 7 Conclusion
We presented a unified theory showing hallucinations are compression failures triggered by insufficient information for rare events. Our framework reconciles the paradox: transformers are “Bayesian in expectation, not in realization” due to positional processing. The QMV theorem provides explicit $O(\log n)$ bounds, EDFL transforms hallucination to quantifiable risk, and operational planners enable prediction and prevention through principled information budgeting. The architectural closure theorem (Theorem 3) shows these theoretical benefits are achievable in practice through ensemble architectures with averaging heads.
@@ Appendix A Appendix A: Proofs and Technical Details
A.2 A.2 Proof of Theorem 2 (Quantified Martingale Violation)
@@
Convexity of $Q\mapsto\operatorname{KL}(P\|Q)$ gives $\mathbb{E}_{\pi}\operatorname{KL}(P\|S_{\pi})\geq\operatorname{KL}(P\|\mathbb{E}_{\pi}S_{\pi})=\operatorname{KL}(P\|\bar{S})$. For the Bernoulli bound, apply data processing to $g:\mathcal{Y}\to\{0,1\}$, $g(y)=\mathbbm{1}_{\{y\in\mathcal{A}\}}$, yielding $\operatorname{KL}(P\|\bar{S})\geq\operatorname{KL}(\operatorname{Ber}(p)\|\operatorname{Ber}(\bar{q}))$. Equality holds iff $g$ is sufficient for distinguishing $P$ from $\bar{S}$. ∎
@@ Appendix F Appendix F: Rare Event Analysis Tables
$\bar{q}$ $\bar{\Delta}$ (nats) $\operatorname{B2T}_{0.05}$ (nats) ISR Decision
@@ Appendix H Appendix H: Factuality Slice Dataset Documentation
H.1 H.1 Dataset Overview
We constructed a custom evidence-grounded QA dataset specifically designed for testing compression-based theories of hallucination. The dataset comprises 3,059 binary classification items with controlled evidence presentation, enabling precise measurement of information sufficiency and permutation sensitivity.
H.2 H.2 Data Sources and Composition
The dataset integrates four established QA resources plus control examples:
FEVER (Fact Verification): 2,000 claims converted to binary true/false questions with Wikipedia evidence from June 2017 snapshot. Each claim paired with gold supporting/refuting evidence sentences and topically-related distractors via BM25 retrieval.
HotpotQA (Multi-hop Reasoning): 2,000 questions requiring evidence synthesis across multiple Wikipedia articles. Preserves multi-hop structure with explicit support spans marking reasoning chains.
NQ-Open (Open-domain QA): 1,000 questions from Natural Questions with evidence retrieved using FEVER Wikipedia as proxy corpus. Answer-bearing sentences marked as support with BM25-selected hard negatives.
PopQA (Long-tail Entities): 500 questions about rare entities (popularity score < 50) testing performance on uncommon knowledge. Evidence retrieved from Wikipedia proxy with weak supervision.
Control Examples: 300+ FEVER "NOT ENOUGH INFO" claims plus synthetic recency traps with outdated evidence, testing model calibration on insufficient/misleading information.
@@
Hard Negative Mining: Up to 120 hard negatives per sample selected via BM25 retrieval, excluding gold evidence. Creates challenging discrimination tasks.
Context Capping: Samples contain up to 60 evidence chunks with all support preserved, then distractors added up to cap. Enables testing at various context lengths n $\in$ [3, 60].
@@
Reproducibility: Complete pipeline code, pinned hashes, and fixed random seeds ensure exact reproducibility. Dataset builder available at: [repository URL withheld for review].

### v2

@@ 7 Conclusion
We connect order sensitivity in evidence-grounded QA to information budgets and a deployment-time abstention interface, yielding an ISR=1{=}1 gate with low hallucination.
@@ Appendix A Appendix A: Proofs and Technical Details
A.2 A.2 Proof of Theorem 2 (Explicit log-scaling under first-order positional sensitivity)
@@
Convexity of $Q\mapsto\operatorname{KL}(P\|Q)$ gives $\mathbb{E}_{\pi}\operatorname{KL}(P\|S_{\pi})\geq\operatorname{KL}(P\|\mathbb{E}_{\pi}S_{\pi})=\operatorname{KL}(P\|\bar{S})$. For the Bernoulli bound, apply data processing to $g:\mathcal{Y}\to\{0,1\}$, $g(y)=\mathbb{1}_{\{y\in\mathcal{A}\}}$, yielding $\operatorname{KL}(P\|\bar{S})\geq\operatorname{KL}(\operatorname{Ber}(p)\|\operatorname{Ber}(\bar{q}))$. Equality holds iff $g$ is sufficient for distinguishing $P$ from $\bar{S}$. ∎
@@ Appendix B Appendix B: Extended Methodology
B.3 B.3 Worked Example: From EDFL to a Decision
Setup. Target hallucination rate $h^{\star}=0.05$ (so $p^{\star}=0.95$). Compute B2T for conservative priors $q_{\text{lo}}$: $\mathrm{B2T}(p^{\star}=0.95;\,q_{\text{lo}})=\mathrm{KL}(\mathrm{Ber}(p^{\star})\,\|\,\mathrm{Ber}(q_{\text{lo}})).$ Numerics (nats): $\begin{array}[]{c|ccc}q_{\text{lo}}&0.02&0.10&0.30\\ \hline\cr\mathrm{B2T}&3.519&1.994&0.963\end{array}$ Budget $\bar{\Delta}$ to reliability. Given $\bar{q}=0.10$, the maximum achievable success at budget $\bar{\Delta}$ solves $\mathrm{KL}(\mathrm{Ber}(p)\,\|\,\mathrm{Ber}(0.10))\leq\bar{\Delta}$. Selected budgets $\to$ $p_{\max}$: $\begin{array}[]{c|cccc}\bar{\Delta}\ (\text{nats})&0.5&1.0&2.0&3.0\\ \hline\cr p_{\max}&0.495&0.689&0.951&\approx 1.000\end{array}$ ISR gate. If $\bar{q}=0.10$ and measured $\bar{\Delta}=2.0$, then $\mathrm{ISR}=\bar{\Delta}/\mathrm{B2T}=2.0/1.994\approx 1.00\Rightarrowanswer$. If instead $q_{\text{lo}}=0.02$, then $\mathrm{B2T}=3.519$ and the same budget gives $\mathrm{ISR}\approx 0.57\Rightarrowabstain$ or acquire info.
@@ Appendix F Appendix F: Rare Event Analysis Tables
$\bar{q}$ $\bar{\Delta}$ (nats) $\operatorname{B2T}_{p^{\star}=0.95}$ (nats) ISR Decision
@@ Appendix H Appendix H: Factuality Slice Dataset Documentation
H.1 H.1 Meta-benchmark Overview
We evaluate on Factuality Slice, a meta-benchmark assembled from five evidence-grounded QA benchmarks: four established benchmarks (FEVER, HotpotQA, NQ-Open, PopQA) plus Controls. We standardize each benchmark into a common evidence-grounded format (question, evidence chunks, support spans, and hard negatives) to enable controlled permutation tests under a shared protocol. The resulting suite contains 3,059 binary adjudication items with controlled evidence presentation, enabling precise measurement of permutation sensitivity and information sufficiency.
Table 7: Factuality Slice: five-benchmark meta-benchmark. Caps are per-source sampling limits before filtering/standardization; the final suite used in this paper contains 3,059 items.
Benchmark Cap Task type Evidence source Why included
FEVER 2,000 Fact verification (T/F) Wikipedia evidence Clean binary adjudication with gold support/refute evidence
HotpotQA 2,000 Multi-hop QA Wikipedia passages Order sensitivity under multi-hop evidence chains
NQ-Open 1,000 Open-domain QA Retrieved Wikipedia Retrieval-style grounding with diverse entities
PopQA 500 Long-tail QA Retrieved Wikipedia Stress-test rare entities (popularity score <50<50)
Controls 300+ NEI + recency Retrieved / synthetic Calibration under insufficient or outdated evidence
H.2 H.2 Benchmarks and Composition
Factuality Slice integrates four established QA benchmarks plus a Controls benchmark:
FEVER (Fact Verification): Up to 2,000 claims converted to binary true/false questions with Wikipedia evidence from a June 2017 snapshot. Each claim paired with gold supporting/refuting evidence sentences and topically-related distractors via BM25 retrieval.
HotpotQA (Multi-hop Reasoning): Up to 2,000 questions requiring evidence synthesis across multiple Wikipedia articles. Preserves multi-hop structure with explicit support spans marking reasoning chains.
NQ-Open (Open-domain QA): Up to 1,000 questions from Natural Questions with evidence retrieved using a Wikipedia proxy corpus. Answer-bearing sentences marked as support with BM25-selected hard negatives.
PopQA (Long-tail Entities): Up to 500 questions about rare entities (popularity score <50<50) testing performance on uncommon knowledge. Evidence retrieved from a Wikipedia proxy with weak supervision.
Controls: 300+ FEVER “NOT ENOUGH INFO” claims plus synthetic recency traps with outdated evidence, testing model calibration on insufficient or misleading information.
Inclusion criteria (brief).
  • FEVER: Include SUPPORTS/REFUTES claims with at least one evidence sentence; map to binary True/False questions. FEVER NEI claims are reserved for Controls.
  • HotpotQA: Include questions with annotated supporting_facts and preserve multi-hop support spans.
  • NQ-Open / PopQA: Include items with non-empty question and answer; retrieve evidence from a Wikipedia proxy corpus and mark answer-bearing sentences as support (weak supervision).
  • PopQA long-tail filter: Restrict to popularity score <50<50.
  • Controls: Include insufficient-evidence items (FEVER NEI) and recency traps with outdated evidence.
@@
Hard Negative Mining: Retrieve up to 120 candidate distractors per sample via BM25, excluding gold evidence; the final context is capped (below).
Context Capping: Cap each sample to at most 60 evidence chunks, preserving all support spans and filling remaining slots with distractors. Enables testing at context lengths $n\in[3,60]$.
@@
Reproducibility: Dataset construction pins hashes and uses fixed random seeds. To preserve double-blind review, code and processed dataset URLs are withheld in the main paper and will be released at camera-ready.